# AI-Assisted Human Evaluation of Machine Translation

### v2

@@ 4.2 ESAAI for Evaluation of WMT Systems
Our goal is for ESAAI to be as reliable or more than ESA in ranking MT systems. We consider MQMWMT collected by Freitag et al. (2023) as the human gold standard and show the system-level correlations with our protocol in App. Figure˜9. Both ESA and ESAAI have similar correlations with MQMWMT, justifying our setup. In Table˜7 we show that this protocol does not stray far away from existing ones in terms of segment-level rating. Many of these cross-protocol correlations are on part with inter-annotator agreement, which is naturally the upper bound. Notably, ESAAI has higher correlation than ESA or MQM by Kocmi et al. (2024) alone.
@@
Table 7: Kendall $\tau_{c}$ segment-level correlations between evaluation protocols, ESA and ESAAI use direct scores.
@@
The goal is to speed-up the annotation process without sacrificing quality. This can be achieved by removing, or automating, redundant decisions and actions on the annotator’s side. In this segment, we do so by skipping high-quality segments for which we can predict that the annotator’s would not mark any error spans.
Our QE system, GEMBA is recall-focused and therefore the occurrence of “false positive” error spans is low. In 89% of cases, spans that were marked by the QE to have 0 errors retained 0 errors after annotation (Table˜8), and such segments had an average score of 95. This makes it possible to also use the QE as a pre-filtering step. If we replace all such segments with 100 (to not overfit), all but one system comparisons remain the same. Alternatively, one can also filter segments for which the QE marks 0 errors for most systems, which has the advantage that we do not alter the data. For this method, again all but one system comparison would be the same (Figure˜7). Pre-filtering can thus result in almost 25% budget saving (~52 segments per system).
@@
Comparing the quality of two annotation protocols is not straightforward because of the absence of a gold standard. We now take a practical perspective where we desire a protocol that finds the true system ordering with as few examples as possible. This does not require any gold standard to compare to. Instead, this approach compares a subset of the annotations of a protocol to the full set of annotations from the same protocol.
Figure 8: Subset consistency accuracy of a system ranking only on a subset against ranking on full data. Percentages are averages across visible subset sizes, which corresponds to the area under the curve. See figure in tabular form in App. Table˜9.
With large enough evaluation, even noisy annotation schemes yield the true system ordering. Conversely, only robust annotation schemes yield this ordering on a small scale. We formalize this in Appendix˜A to show that ESAAI leads to better annotations than ESA or MQM. We measure the accuracy of the ordering ($m_{1}{>_{I}}m_{2}$) of systems ($\mathcal{M}$) computed on a subset of segments ($I$) against the ordering given by the full data ($a_{m_{1}}{>}a_{m_{2}}$):
@@
This simulates the setup where we wish to save costs with fewer number of annotations but care about arriving at the true system ordering. We refer to this metric as subset consistency accuracy and in Appendix˜A justify that annotation protocols with lower noise reach higher accuracies. For each subset size, e.g. 30 source sentences, we select 1000 random subsets and compute the system ranking accuracy.
Results in Figure˜8 suggest that the QE system, GEMBA, alone is the most consistent. However, this is solely because it can be perceived as a single annotator and thus there is no inter-annotator confusion. From human evaluation protocols, ESAAI with scores has the highest stability and quality of scores. This is not the result of automation bias because only the annotators themselves assign the final score. In practice this would mean that one can annotate fewer examples (e.g. 2000 for ESAAI) to obtain the same system-level accuracy as lower-quality protocol (e.g. 2500 for ESA), thus further lowering the costs.
@@ 5 Conclusion
Our AI-assisted protocol of human evaluation of MT is faster and cheaper. This protocol is more robust and self-consistent and increases inter-annotator agreement by priming the annotators with pre-annotated error spans. Our analysis also shows that the annotators did not over-rely on the AI and were able to maintain evaluation quality. The inclusion of AI in evaluation also opens many options for further evaluation economy by reducing the test set size requirements. To this end, we introduce subset consistency accuracy, which quantifies how many annotations could be saved while arriving at similar final system ordering.
@@
Despite the advantages in lower costs per error span of the presented setup, we urge practitioners to not use this approach when metrics evaluation is one of the expected tasks due to the particular bias to the used metric in the setup. The intended application of this pipeline is purely a more efficient evaluation of machine translation system quality.
Both ESAAI and GEMBA rank GPT-4-5shot as the best system, a system that uses the same LLM to translate sentences as we use to generate for GEMBA. This indicated a weakness that our approach is biased towards systems build on top of the same underlying LLM. Liu et al. (2023) described this phenomena when the same system used for generating output should not be used to also evaluate them. This issue could be mitigated by using two different LLMs to generate error spans.
@@
The annotators were paid a standard commercial translator wage in the respective country. No personal data was collected and the showed data was screened for potentially disturbing content.
@@ Appendix A Subset Consistency Formalization
This section justifies the setup in Section˜4.2 and is reminiscent of the work of Riley et al. (2024). A key distinction is that we are considering ranking stability with respect to the protocol itself. We do so by bootstrapping subsets of the data.
@@ Appendix B User Guidelines
The following are are annotation guidelines for our two local ESAAI campaigns, which is closely based on the setup of Kocmi et al. (2024).

### v3

@@ 4.2 ESAAI for Evaluation of WMT Systems
Our goal is for ESAAI to be as reliable as or more reliable than ESA in ranking MT systems. We consider MQMWMT collected by Freitag et al. (2023) as the human gold standard and show the system-level correlations with our protocol in Appendix Figure˜9. Both ESA and ESAAI have similar correlations with MQMWMT, justifying our setup. In Table˜7 we show that this protocol does not stray far from existing ones in terms of segment-level rating. Many of these cross-protocol correlations are on part with inter-annotator agreement, which is naturally the upper bound. In particular, ESAAI has a higher correlation than ESA or MQM by Kocmi et al. (2024b) alone.
@@
Table 7: Kendall $\tau_{c}$ segment-level correlations between evaluation protocols. ESA and ESAAI use direct scores.
@@
The goal is to speed up the annotation process without sacrificing quality. This can be achieved by removing, or automating, redundant decisions and actions on the annotator’s side. In this segment, we do so by skipping high-quality translations for which we can predict that the annotator’s would not mark any error spans.
Our QE system, GEMBA, is recall-focused, and therefore the occurrence of “false positive” error spans is low. In 89% of the cases, the QE marked the spans as having 0 errors and retained 0 errors after the annotation (first row in Table˜8), and these segments have an average score of 95. This makes it possible to also use the QE as a prefiltering step. If we replace all such segments with 100 (not to overfit), all but one system comparisons remain the same (Figure˜7, left). Alternatively, one can also exclude segments for which the QE marks 0 errors for most systems, which has the advantage that we do not alter the data. For this method, again all but one system comparison would be the same (Figure˜7, right). Pre-filtering can thus result in almost 25% budget saving (~52 segments per system).
@@
Comparing the quality of two annotation protocols is not straightforward because of the absence of a gold standard. We now take a practical perspective where we desire a protocol that finds the true system ordering with as few examples as possible. This does not require any gold standard to compare with. Instead, this approach compares a subset of the annotations of a protocol with the full set of annotations from the same protocol.
Figure 8: Subset consistency accuracy of a system ranking only on a subset against ranking on full data. Percentages are averages across visible subset sizes, which corresponds to the area under the curve. See figure in tabular form in Appendix Table˜9.
With a sufficiently large evaluation, even noisy annotation schemes yield the true system ordering. Conversely, only robust annotation schemes yield this ordering on a small scale. We formalize this in Appendix˜A to show that ESAAI leads to better annotations than ESA or MQM. We measure the accuracy of the ordering ($m_{1}{>_{I}}m_{2}$) of systems ($\mathcal{M}$) computed on a subset of segments ($I$) against the ordering given by the full data ($a_{m_{1}}{>}a_{m_{2}}$):
@@
This simulates the setup where we wish to save costs with fewer number of annotations but care about arriving at the true system ordering. We refer to this metric as subset consistency accuracy, similar to split-half reliability, and in Appendix˜A justify that annotation protocols with lower noise reach higher accuracies. For each subset size, e.g. 30 source sentences, we select 1000 random subsets and compute the system ranking accuracy.
The results in Figure˜8 suggest that the QE system, GEMBA, alone very consistent. However, this is solely because it can be perceived as a single annotator and thus there is no inter-annotator confusion. Overall and among human evaluation protocols, ESAAI with scores has the highest stability and quality of scores. This is not the result of automation bias because only the annotators themselves assign the final score. In practice, this would mean that one can annotate fewer examples (e.g. 2000 for ESAAI) to obtain the same system-level accuracy as lower-quality protocol (e.g. 2500 for ESA), thus further lowering the costs.
@@ 5 Conclusion
Our AI-assisted protocol of human evaluation of MT is faster and cheaper. This protocol is more robust and self-consistent and increases inter-annotator agreement by priming the annotators with pre-annotated error spans. Our analysis also shows that the annotators did not overrely on the AI and were able to maintain evaluation quality. The inclusion of AI in evaluation also opens many options for further evaluation economy by reducing the test set size requirements. To this end, we introduce subset consistency accuracy, which quantifies how many annotations could be saved while arriving at a similar final system ordering.
@@
Despite the advantages in lower costs per error span of the presented setup, we urge practitioners not to use this approach when metrics evaluation is one of the expected tasks due to the particular bias to the used metric in the setup. The intended application of this pipeline is purely a more efficient evaluation of the quality of the machine translation system.
Both ESAAI and GEMBA rank GPT-4-5shot as the best system, a system that uses the same LLM to translate sentences as we use to generate for GEMBA. This indicated a weakness that our approach is biased towards systems built on top of the same underlying LLM. Liu et al. (2023) described this phenomenon when the same system used to generate output should not be used to also evaluate them. This issue could be mitigated by using two different LLMs to generate error spans.
@@
The annotators were paid a standard commercial translator wage in the respective country. No personal data was collected and the data shown to the annotators was screened for potentially disturbing content.
@@ Appendix A Subset Consistency Formalization
This section justifies the setup in Section˜4.2 and is reminiscent of the work of Riley et al. (2024) or split-half reliability. A key distinction is that we are considering ranking stability with respect to the protocol itself. We do so by bootstrapping subsets of the data.
@@ Appendix B User Guidelines
The following are are annotation guidelines for our two local ESAAI campaigns, which is closely based on the setup of Kocmi et al. (2024b).